OpenStudy (anonymous):

What is the simplified form of the expression –(10)^–1?

OpenStudy (anonymous):

Easy... What is -10 minus 1?

OpenStudy (anonymous):

Ohhh wait I see, hold on.

OpenStudy (anonymous):

I may or may not be correct... but I think what you do is.. For every negative power of a number, you move the decimal that many units to the left. So if -10.0 is to the power of negative 1, then I believe the simplified answer may be -1. Again, I'm not quite sure, but I think that should be right. If anyone knows better then please correct me.

OpenStudy (anonymous):

the answer choices are -1/10, \[-\frac{ 1}{ 10 ^10 },-\frac{ 1 }{ 10}, \], 1/10, 10

OpenStudy (anonymous):

-(10)^-1=-1/10